# A Multimodal Data Fusion Generative Adversarial Network for Real Time Underwater Sound Speed Field Construction

### v1

@@ Abstract
Sound speed profiles (SSPs) are essential parameters underwater that affects the propagation mode of underwater signals and has a critical impact on the energy efficiency of underwater acoustic communication and accuracy of underwater acoustic positioning. Traditionally, SSPs can be obtained by matching field processing (MFP), compressive sensing (CS), and deep learning (DL) methods. However, existing methods mainly rely on on-site underwater sonar observation data, which put forward strict requirements on the deployment of sonar observation systems. To achieve high-precision estimation of sound velocity distribution in a given sea area without on-site underwater data measurement, we propose a multi-modal data-fusion generative adversarial network model with residual attention block (MDF-RAGAN) for SSP construction. To improve the model’s ability for capturing global spatial feature correlations, we embedded the attention mechanisms, and use residual modules for deeply capturing small disturbances in the deep ocean sound velocity distribution caused by changes of SST. Experimental results on real open dataset show that the proposed model outperforms other state-of-the-art methods, which achieves an accuracy with an error of less than 0.3m/s. Specifically, MDF-RAGAN not only outperforms convolutional neural network (CNN) and spatial interpolation (SITP) by nearly a factor of two, but also achieves about 65.8% root mean square error (RMSE) reduction compared to mean profile, which fully reflects the enhancement of overall profile matching by multi-source fusion and cross-modal attention.
keywords:
Sound speed profile (SSP) , Sound speed field construction , Generative adversarial network (GAN) , Data fusion
journal: Knowledge-Based Systems
[…]
{highlights}
A MDF-RAGAN model that quickly obtain sound velocity distribution without the need for underwater on-site data measurement.
A residual block that makes the model learn the influence of SST on the variation of sound velocity.
Embedded attention mechanism that makes the model learn the correlation of sound velocity distribution in large-scale spatial regions.
1 Introduction
The estimation of underwater sound velocity distribution has become a research hotspot in recent years, as it affects the propagation mode of underwater signals and therefore has a critical impact on the energy efficiency of underwater acoustic communication and accuracy of underwater acoustic positioning (Erol-Kantarci et al., 2011; Piccolo et al., 2019; Luo et al., 2021).
@@ 1 Introduction
Traditionally, sound speed profiles (SSPs), that are popular adopted to describe the distribution of sound speed, are usually measured by conductivity, temperature, and depth profiler (CTD) or sound velocity profiler (SVP) (Wang et al., 2014), but these instruments rely on shipborne deployment systems, leading to expensive economic and labor costs. To improve the efficiency of measuring sound velocity distribution, scholars have proposed several methods for inverting sound velocity distribution leveraging sonar observation data, based on the influence of sound velocity distribution on the sound field. These methods can be mainly divided into 3 categories: matching field processing (MFP) (Tolstoy et al., 1991; Li and Zhang, 2010; Zhang et al., 2012, 2015), compressive sensing (CS) (Bianco and Gerstoft, 2017; Choo and Seong, 2018), and deep learning (DL) (Stephan et al., 1995; Huang et al., 2021b, 2023).
MFP is a method of determining the sound velocity distribution by simulating the SSP and its corresponding theoretical sound field distribution, and matching it with the measured sound field based on ray theory or normal wave theory. In order to improve the computational efficiency of MFP, heuristic algorithms, such as particle swarm optimization (PSO) (Tolstoy et al., 1991) and genetic algorithm (GA) (Zhang et al., 2015), are usually introduced for accelerating the search process of feature vector coefficients. However, heuristic algorithms still have very high computational complexity and the results obtained are prone to getting stuck in local optima, which affects the accuracy of sound velocity inversion.
@@
Considering that ray theory or normal wave theory can only establish a mapping relationship from sound velocity distribution to sound field distribution, scholars have begun to explore how to directly establish a mapping relationship from sound field distribution to sound velocity distribution. The CS technology achieves this process through a feature mapping dictionary, which further improves the efficiency of sound speed estimation compared to MFP. However, the dictionary establishment process introduces linearization approximation (Bianco and Gerstoft, 2017), which sacrifices the accuracy of sound speed construction.
@@
During the last decades, DL has obtained significant achievement in the field of data processing, benefiting from its ability to handle complex nonlinear mapping relationships. The DL based sound velocity inversion method not only reduces accuracy loss, but also maintains the real-time advantage at the working stage (training can be completed in advance) (Huang et al., 2021b, 2023). However, above mentioned SSP inversion methods highly rely on sonar observation data, which put forward strict requirements on the deployment of sonar observation systems, making it difficult to achieve large-scale ocean coverage.
@@
Recently, to eliminate the need for underwater on-site data measurement, scholars have proposed two ways of approaches to improve the efficiency of sound velocity estimation: one is using historical sound velocity data to forecast the current sound velocity from the perspective of time series prediction, and the other is fusing the remote sensing sea surface temperature (SST) data and historical sound velocity data to estimate the sound velocity distribution from the perspective of spatial sound velocity field construction. In 2023, Piao et al. proposed an SSP prediction method under internal waves based on long short-term memory (LSTM) neural networks (Piao et al., 2023). Lu et al. further developed a hierarchical LSTM (H-LSTM) model that predict sound speed distribution at different depth layers (Lu et al., Jan., 2024). However, the sound velocity prediction method is only applicable to areas with sufficient historical sound velocity distribution, which can not be satisfied in some ocean areas due to difficulty in measuring SSPs.
@@
To tackle this issue, Li et al. (Li et al., Aug., 2021) and Wu et al. (Wu et al., 2024) both proposed SSP construction methods fusing SST data and historical SSPs, where a self-organized mapping (SOM) neural network was proposed in (Li et al., Aug., 2021) and a convolutional neural network (CNN) was proposed in (Wu et al., 2024). However, both (Li et al., Aug., 2021) and (Wu et al., 2024) focus excessively on local spatial correlations while neglecting global spatial relationships, resulting in substantial accuracy limitations.
@@
To achieve high-precision estimation of sound velocity distribution in a given sea area without on-site underwater data measurement, we propose a multi-modal data-fusion generative adversarial network model with residual attention block (MDF-RAGAN) for SSP construction. We aim to improve the model’s ability for capturing global spatial feature correlations by the attention mechanisms, and use residual modules for deeply capturing small disturbances in the deep ocean sound velocity distribution caused by changes of SST. The contributions of this work are summarized as follows.
  • 1.
    To quickly obtain sound velocity distribution of a given area, we proposed a MDF-RAGAN model that fusing SST data with historical sound velocity distribution, which eliminates the need for underwater on-site data measurement. The generator is used to establish a mapping relationship from the fused data to the sound velocity distribution, while the discriminator is used to constrain the trend of the sound velocity distribution.
  • 2.
    To improve the accuracy of underwater sound velocity construction, we integrated the residual attention block into the multi-modal data-fusion GAN model. With the residual block, the model could learn the influence of SST on the variation of sound velocity. With the attention mechanism, the model could learn the correlation of sound velocity distribution in large-scale spatial regions.
  • 3.
@@
The rest part of this paper is organized as follows. Section 2 describes the MDF-RAGAN in detail. Section 3 evaluates and discusses the performance of MDF-RAGAN on real-word open dataset. The conclusion is given in Section 4.
2 Sound Speed Field Construction by MDF-RAGAN
2.1 Overview of MDF-RAGAN Framework
In this paper, we aim to construct the SSP of a target location with the help of SST data and SSP distribution of neighboring locations. This task is significantly challenging: firstly, the spatial coordinates, SST and SSPs are heterogeneous modal data, which need to be efficiently aligned and fused in a unified feature space; secondly, the SSPs show highly nonlinear variations with environmental variables such as depth, temperature, salinity, etc., which puts stringent requirements on the network’s expressive ability.
@@ 2.1 Overview of MDF-RAGAN Framework
To address these challenges, we propose the MDF-RAGAN, a novel framework consisting of three key components as shown in Fig. 1. We first map three types of heterogeneous data modalities, namely, geographic coordinates of the target point and its surrounding grid, SST, and historical SSPs, into a unified feature space by multi-source data fusion preprocessing; subsequently, we use the cross-modal perturbation attention block to capture the spatial correlation between sound velocity distributions; finally, the end-to-end optimization strategy is constructed by combining the generative adversarial training and the multi-task regression loss to ensure a high level of prediction performance under a small number of samples.
The input of the model are coordinates, SST and historical SSPs. Let the set of SSPs at the reference points surrounding a given latitude-longitude coordinate be denoted as $\mathbf{R}_{ssp}^{n}\in\mathbb{R}^{N\times D}$, where $N$ denotes the number of surrounding reference points, and $D$ represents the depth dimension. The geographic coordinates (longitude $\lambda$, latitude $\phi$) and the SST obtained by remote sensing at each reference point are denoted by $\mathbf{R}_{loc}^{n}\in\mathbb{R}^{N\times 2},\quad\mathbf{R}_{sst}^{n}\in\mathbb{R}^{N\times 1}$. The geographic coordinates and the sea surface temperature at the target point are denoted as $\mathbf{R}_{loc}^{t}\in\mathbb{R}^{1\times 2},\mathbf{R}_{sst}^{t}\in\mathbb{R}^{1\times 1}$.
The final output from the generator network is a small perturbation $\mathbf{\delta}_{ssp}$ at the target location. Therefore, the estimated SSP $\mathbf{\hat{s}}_{ssp}$ at the target point can be expressed as the sum of the mean SSP of surrounding reference points and the generated perturbation:
@@
This decomposition approach guides the generator network to focus explicitly on the local variability at the target point. Leveraging the overall trend information from surrounding reference points, we represent the estimated SSP as the sum of a mean profile of surrounding references and a small perturbation term. The perturbation term is learned by the generator network to compensate for inadequacies inherent in the mean profile model, especially under complex marine conditions, thereby effectively enhancing prediction accuracy. Moreover, such decomposition simplifies model training, enabling the network to focus more effectively on capturing deep nonlinear deviations rather than merely reconstructing the background mean.
The detailed model of MDF-RAGAN is given in Fig. 2. In the following section, we will introduce the generator, discriminator, and optimization strategy step by step.
Figure 2: The proposed MDF-RAGAN model for SSP estimation, which consists of a generator and a discriminator, incorporating FMB, LFB, and CMPAB for effective multi-source data fusion.
2.2 Generator
The core objective of the generator is to map the fused features from multiple sources to detailed perturbations of the SSP at the target location, and thus reconstruct temperature-induced variations in the ocean on the sound velocity distribution. The entire network consists of three stages in sequence: feature encoding and fusion, cross-modal cross-attention, and perturbation decoding.
2.2.1 Multi-Source Data Fusion Preprocessing
The SSP shows complex nonlinear behaviors with the variations in ocean depth, temperature and salinity. The geographic coordinates identify the thermohaline structure of the ocean region, while the SST distribution directly affects the surface thermohaline gradient, which in turn has a significant effect on the upper sound speed distribution. To accurately estimate the SSP at target points, it is necessary to effectively align and fuse multi-sources of data, including geographic coordinates and SST, within a unified feature space to make full use of their physical complementarities.
@@ 2.2.1 Multi-Source Data Fusion Preprocessing
The label fusion block (LFB) aims to adaptively fuse the label information from both the target point and surrounding reference points. Specifically, the latitude and longitude coordinates and SST obtained by remote sensing are first embedded into the feature space $d_{r}$ through their corresponding encodings, each undergoing linear mapping followed by layer normalization:
@@
where $W_{loc}\in\mathbb{R}^{2\times d_{r}}$ and $W_{sst}\in\mathbb{R}^{1\times d_{r}}$ are the linear transformation matrices that map these two types of inputs to the $d_{r}$-dimensional feature space, with the corresponding bias terms being $b_{loc}\in\mathbb{R}^{d_{r}}$ and $b_{sst}\in\mathbb{R}^{d_{r}}$. Thus, we can obtain the $d_{r}$-dimensional embedding feature vectors for surface remote sensing temperature and coordinate points.
The same feature transformations are applied to the corresponding reference point data, mapping the reference point features $\left\{\mathbf{z}_{sst}^{r,n},\mathbf{z}_{loc}^{r,n}\right\}_{n}^{N}$ into a consistent dimension. The aligned labels are concatenated along the sample dimension to form the fusion matrix $\mathbf{Z}^{r}\in\mathbb{R}^{N\times d_{r}}$. Subsequently, the LFB is employed to fuse the features across both the sample and feature dimensions. An initial fusion is performed via a learnable convolutional transformation:
$\mathbf{\tilde{Z}^{r}}=\gamma_{1}\frac{\mathbf{Z}^{r}-\mu_{Z}}{\sigma_{Z}}+\beta_{1},$ (4)
$\mathbf{\bar{Z}}^{r}=\gamma_{2}\frac{\mathbf{W}\times\mathbf{\tilde{Z}^{r}}+\mathbf{b}-\mu_{Y}}{\sigma_{Y}}+\beta_{2},$ (5)
where $\mu_{Z},\sigma_{Z}$ are the mean and standard deviation computed along the channel dimension of the matrix $\mathbf{Z}^{r}$, $\mu_{Y},\sigma_{Y}$ denote the corresponding statistics computed after the convolution operation, and $\gamma_{1},\beta_{1}$ and $\gamma_{2},\beta_{2}$ are the learnable scaling and shifting parameters normalized twice. After the transformation, we divide the fusion matrix $\mathbf{\bar{Z}}^{r}$ into two parts $\mathbf{\bar{Z}}_{A}^{r},\mathbf{\bar{Z}}_{B}^{r}\in\mathbb{R}^{N\times d_{r}}$ and feed them to the linear gating unit GLU:
@@
This enables adaptive weighting across labels from different modalities. The proposed mechanism dynamically adjusts the relative importance of coordinates and temperature in SSP predictions according to the geographic and thermal conditions of various ocean regions.
@@
Considering that the detailed variations of SSP along the depth dimension are closely associated with the locations of scattering layers and thermoclines, the Feature Mapping Block (FMB) further enhances these physical characteristics through local context modeling. Given the SSP features at the reference points $\mathbf{Z}_{ssp}^{r}\in\mathbb{R}^{1\times N\times d_{r}}$, the FMB initially extracts fine-grained, depth-wise features through depthwise separable convolutions:
@@
Specifically, the calculation can be expressed as follows:
$Y_{i,j}=\sum_{u=1}^{k_{h}}{\sum_{v=1}^{k_{w}}{K_{u,v}^{\left(dw\right)}\left(\mathbf{\dot{Z}}_{ssp}^{r}\right)_{i+u-1,j+v-1}}},$ (8)
$\left(\mathbf{\dot{Z}}_{ssp}^{r}\right)_{i,j}=K_{d}^{\left(pw\right)}Y_{i,j}+b_{d},$ (9)
where $K^{\left(dw\right)}\in\mathbb{R}^{1\times k_{h}\times k_{w}}$ is the depthwise convolution kernel, which performs $k_{h}\times k_{w}$ convolutions independently on each channel. $K^{\left(pw\right)}\in\mathbb{R}^{1\times 1\times 1\times 1}$ and $b\in\mathbb{R}^{1}$ denote the kernel and bias used in the subsequent pointwise convolution. In deep-sea environments, abrupt variations in salinity across adjacent depth levels may amplify gradients during training. To mitigate potential gradient explosion introduced by convolutional operations alone, we further apply a SimpleGate activation module after convolution:
@@
where the weight $\xi$ is a learnable scaling factor used to balance the relative contributions of the original depth features $\mathbf{Z}_{ssp}^{r}$ and the enhanced features $SG\left(\mathbf{\dot{Z}}_{ssp}^{r}\right)$ in the residual link. Through training, the model automatically adjusts the size of $\xi$: when $\xi$ is large, it emphasises the local nonlinear changes captured by the depthwise separable convolution and SimpleGate; when $\xi$ is small or even close to 0, it focuses more on preserving the overall trend of the SSP.
@@
Through synergistic interaction between the LFB and FMB, the module achieves the unified alignment and enhancement of multimodal labeling information with deep spatial context features. This integration provides MDF-RAGAN with rich, structured, and physically meaningful initial features, thereby supporting robust and precise adversarial learning.
2.2.2 Cross-Modal Perturbation Attention Block

### v3

@@ Abstract
Sound speed profiles (SSPs) are crucial underwater parameters that determine the propagation patterns of acoustic signals, directly influencing the energy efficiency of underwater communication and the accuracy of positioning systems. Conventional techniques for obtaining SSPs, such as matched field processing (MFP), compressive sensing (CS), and deep learning (DL), typically depend on in-situ sonar measurements, which impose stringent requirements on the deployment of underwater observation systems. To overcome this limitation and enable high-precision sound speed field reconstruction without the need for on-site underwater data collection, we propose a novel multimodal data-fusion generative adversarial network enhanced with residual attention blocks (MDF-RAGAN). This architecture integrates attention mechanisms to capture global spatial feature correlations effectively, while residual modules are employed to extract subtle perturbations in deep-ocean sound velocity distribution caused by sea surface temperature (SST) variations. Experimental results on a public real-world dataset demonstrate that the proposed model outperforms other state-of-the-art methods, achieving an estimating error of less than 0.3 m/s. Specifically, MDF-RAGAN reduces the root mean square error (RMSE) by nearly half compared to convolutional neural network (CNN) and spatial interpolation (SITP) methods, and attains a 65.8% RMSE reduction relative to the mean profile method. These results highlight the effectiveness of multi-source fusion and cross-modal attention in enhancing the accuracy and robustness of sound speed profile reconstruction.
I Introduction
The estimation of underwater sound speed profiles (SSPs) has emerged as a significant research focus, owing to its critical role in governing acoustic signal propagation. This, in turn, directly determines the energy efficiency of underwater acoustic communication systems and the precision of underwater positioning [1, 2, 3, 4, 5].
@@ I Introduction
Sound speed profiles (SSPs) are popular adopted to describe the distribution of underwater sound speed [6], and they are traditionally measured by conductivity, temperature, and depth profiler (CTD) [7, 8, 9] or sound velocity profiler (SVP) [10]. The main drawback of instrument measurement methods is their low spatiotemporal resolution. To enhance the efficiency of measuring sound velocity profiles, researchers have developed various inversion methods that utilize sonar observation data by exploiting the dependence of acoustic field properties on the sound speed distribution. These methods can be mainly divided into 3 categories: matching field processing (MFP) [11, 12, 13, 14], compressive sensing (CS) [15, 16], and deep learning (DL) [17, 18, 19, 20].
MFP is a method of determining the sound velocity distribution by simulating the SSP and its corresponding theoretical sound field distribution, and matching it with the measured sound field based on ray theory or normal wave theory. However, the MFP method has high algorithm complexity, which reduces its execution efficiency and makes it difficult to continuously construct the sound velocity field.
@@
Considering that ray theory or normal wave theory can only establish a mapping relationship from sound velocity distribution to sound field distribution, scholars have begun to explore how to directly establish a mapping relationship from sound field distribution to sound velocity distribution. The CS technology achieves this process through a feature mapping dictionary, which further improves the efficiency of sound speed estimation compared to MFP. However, the dictionary establishment process introduces linearization approximation [15], which sacrifices the accuracy of sound speed construction.
@@
During the last decades, DL has obtained significant achievement in the field of data processing, benefiting from its ability to handle complex nonlinear mapping relationships. The DL based sound velocity inversion method not only reduces accuracy loss, but also maintains the real-time advantage at the working stage (training can be completed in advance) [18, 19]. However, the aforementioned sound speed profile (SSP) inversion methods are highly dependent on sonar observation data. This imposes stringent requirements on the deployment of sonar observation systems, thereby hindering their capability for large-scale oceanographic coverage.
@@
Recently, to avoid the need for in-situ underwater measurements, researchers have introduced two alternative methods for efficient sound speed estimation. The first approach employs historical sound velocity data to predict current sound speed profiles using time-series forecasting techniques. The second method integrates remotely sensed sea surface temperature (SST) data with historical sound velocity records to reconstruct the sound speed field through spatial interpolation and data fusion techniques. In 2023, Piao et al. proposed an SSP prediction method under internal waves based on long short-term memory (LSTM) neural networks [21]. Lu et al. further developed a hierarchical LSTM (H-LSTM) model that predict sound speed distribution at different depth layers [22]. However, the sound velocity prediction method is only applicable to areas with sufficient historical sound velocity distribution, which can not be satisfied in some ocean areas due to difficulty in measuring SSPs.
@@
To solve this problem, Li et al. [23] and Wu et al. [24] proposed SSP construction methods fusing SST data and historical SSPs, where a self-organized mapping (SOM) neural network was proposed in [23] and a convolutional neural network (CNN) was proposed in [24]. However, the approaches proposed by [23] and [24] are primarily confined to modeling adjacent spatial dependencies. Their failure to capture long-range global relationships thus imposes significant limitations on model accuracy.
@@
To achieve high-precision estimation of the sound speed profile (SSP) in a given sea area without relying on in situ measurements, we propose a multi-modal data-fusion generative adversarial network incorporating a residual attention block (MDF-RAGAN). This model is designed to enhance the capture of global spatial feature correlations through attention mechanisms, while residual modules are employed to extract fine-grained perturbations in deep-ocean sound velocity distribution caused by SST variations. The contributions of our work are concluded as follows:
  • To efficiently obtain the sound velocity distribution in a given ocean area, we propose an MDF-RAGAN model that integrates SST data with historical SSPs, thereby eliminating the reliance on in-situ underwater measurements. In this framework, the generator learns a mapping from the fused multimodal inputs to the target sound speed distribution, while the discriminator constrains the spatiotemporal consistency and physical plausibility of the generated profiles.
  • To enhance the accuracy of underwater SSP reconstruction, we introduce a residual attention mechanism into a multi-modal data-fusion GAN. The residual blocks enable the model to capture subtle variations in sound velocity caused by SST changes, while the attention module improves its ability to model long-range spatial dependencies and regional correlations within large-scale oceanic environments.
@@
The rest part of this paper is organized as follows. Section II provides a detailed description of the proposed MDF-RAGAN framework. In Section III, the model is thoroughly evaluated and its performance is discussed based on experiments conducted on a real-world open dataset. Finally, Section IV concludes the paper and outlines potential directions for future research.
II Sound Speed Estimation by MDF-RAGAN
II-A Overview of MDF-RAGAN Framework
In this study, we seek to reconstruct the SSP at a target location using remote sensing SST data and the historical SSP distributions. This task presents two major challenges: first, the spatial coordinates, SST values, and SSP measurements constitute heterogeneous modal data, requiring effective alignment and fusion within a unified feature representation; second, the SSPs exhibit highly nonlinear dependencies on environmental factors such as depth, temperature, and salinity, demanding a model with strong expressive power to capture these complex relationships.
@@ II-A Overview of MDF-RAGAN Framework
To address these challenges, we propose MDF-RAGAN, a novel framework comprising three key components, as illustrated in Fig. 1. First, we map three types of heterogeneous input modalities, namely, the geographic coordinates of the target point and its surrounding grid, SST values, and the average profile of historical SSPs, into a unified feature space through multi-source data fusion preprocessing. Subsequently, a cross-modal perturbation attention mechanism is employed to capture spatial correlations within sound velocity distributions. Finally, an end-to-end optimization strategy is established by integrating generative adversarial training with multi-task regression loss, ensuring robust prediction performance even with limited samples.
In physical oceanography, underwater sound-speed profiles are governed by temperature, salinity, and pressure, with temperature exerting the dominant control in the upper ocean. As the most accessible and spatially continuous thermal observable, sea surface temperature (SST) provides a key boundary condition on mixed-layer temperature and upper-ocean stratification, and previous studies have shown that basin-scale sound-speed structures can be reconstructed from sea-surface variables such as SST and sea surface height (SSH), underscoring the tight coupling between surface thermodynamics and subsurface acoustics [25].
Under climatological and regional regimes, the vertical temperature profile $T(z)$ therefore exhibits a statistically stable dependence on SST, geographic location, and season, which has been extensively exploited in Argo-based empirical reconstructions and oceanographic climatologies. Building on this physical prior, MDF-RAGAN integrates SST, historical SSPs, and geographic coordinates so that SST constrains the surface and mixed-layer state, historical SSPs encode the large-scale background structure, and their joint representation guides the reconstruction of full-depth SSPs in data-sparse regions.
The input of the model are coordinates, SST and historical SSPs. Let the set of SSPs at the reference points surrounding a given latitude-longitude coordinate be denoted as $\mathbf{R}_{ssp}^{n}\in\mathbb{R}^{N\times D}$, where $N$ denotes the number of surrounding reference points, and $D$ represents the depth dimension. The geographic coordinates (longitude $\lambda$, latitude $\phi$) and the SST obtained by remote sensing at each reference point are denoted by $\mathbf{R}_{loc}^{n}\in\mathbb{R}^{N\times 2},\quad\mathbf{R}_{sst}^{n}\in\mathbb{R}^{N\times 1}$. The geographic coordinates and the SST at the target point are denoted as $\mathbf{R}_{loc}^{t}\in\mathbb{R}^{1\times 2},\mathbf{R}_{sst}^{t}\in\mathbb{R}^{1\times 1}$.
The final output from the generator network is a small perturbation $\mathbf{\delta}_{ssp}$ at the target location. Therefore, the estimated SSP $\mathbf{\hat{s}}_{ssp}$ at the target point can be expressed as the sum of the average SSP of surrounding reference points and the generated perturbation:
@@
This decomposition strategy guides the generator network to explicitly focus on capturing variations at the target location. By leveraging the general trend information derived from surrounding reference points, the estimated SSP is expressed as the sum of the average profile of nearby references and a residual perturbation term. The generator is designed to learn this perturbation term, which compensates for the limitations of the mean profile, particularly under complex marine environmental conditions, thereby significantly improving prediction accuracy. Furthermore, this decomposition simplifies the training process, allowing the network to concentrate on modeling deep nonlinear deviations rather than reconstructing the background mean field.
Figure 2: The proposed MDF-RAGAN model for SSP estimation.
The detailed model of MDF-RAGAN is given in Fig. 2. In the following section, we detail the architectures of the generator and discriminator, along with the optimization strategy.
II-B Generator
The primary objective of the generator is to map fused multi-source features into detailed SSP perturbations at the target location, thereby reconstructing ocean temperature-induced variations in the sound velocity distribution. The entire network operates sequentially through three stages: feature encoding and fusion, cross-modal cross-attention, and perturbation decoding.
II-B1 Multi-Source Data Fusion
The SSP exhibits complex nonlinear behavior due to variations in ocean depth, temperature, and salinity. Geographic coordinates help identify the thermohaline structure of a given ocean region, while SST distribution directly influences the surface thermohaline gradient, significantly affecting the upper ocean sound speed distribution. To accurately estimate the SSP at target points, it is essential to effectively align and fuse multi-source data, such as geographic coordinates and SST, within a unified feature space, thereby leveraging their complementary physical relationships.
@@ II-B1 Multi-Source Data Fusion
The label fusion block (LFB) is designed to adaptively integrate label information from both the target point and its surrounding reference points. Specifically, the latitude-longitude coordinates and remote sensed SST data are first embedded into a feature space of dimension $d_{r}$ via respective encoding pathways. Each feature stream is processed through a linear transformation followed by layer normalization:
@@
where $W_{loc}\in\mathbb{R}^{2\times d_{r}}$ and $W_{sst}\in\mathbb{R}^{1\times d_{r}}$ are the linear transformation matrices that map these two types of inputs to the $d_{r}$-dimensional feature space, with the corresponding bias terms being $b_{loc}\in\mathbb{R}^{d_{r}}$ and $b_{sst}\in\mathbb{R}^{d_{r}}$. Thus, we can obtain the $d_{r}$ dimensional embedding feature vectors for surface remote sensing temperature and coordinate points.
The same feature transformations are applied to the corresponding reference point data, mapping the reference point features into a consistent dimension. The aligned labels are concatenated along the sample dimension to form the fusion matrix $\mathbf{Z}^{r}\in\mathbb{R}^{N\times d_{r}}$, which is then processed through a learnable fusion transformation with instance normalization and learnable scaling/shifting parameters. After the transformation, we divide the fusion matrix $\mathbf{\bar{Z}}^{r}$ into two parts $\mathbf{\bar{Z}}_{A}^{r},\mathbf{\bar{Z}}_{B}^{r}\in\mathbb{R}^{N\times d_{r}}$ and feed them to the linear gating unit GLU:
@@
This enables adaptive weighting across multi-modal labels by dynamically adjusting the relative importance of coordinates and temperature in SSP predictions based on the geographic and thermal conditions of different ocean regions.
@@
Since detailed SSPs vary with depth due to scattering layers and thermoclines, the feature mapping block (FMB) employs local context modeling to amplify these critical physical features. Given the SSP features at the reference points $\mathbf{Z}_{ssp}^{r}\in\mathbb{R}^{1\times N\times d_{r}}$, the FMB initially extracts fine-grained, depth-wise features through depthwise separable convolutions:
@@
The depthwise separable convolution first applies depthwise convolution to each channel independently, followed by a pointwise convolution to fuse channel information. In deep-sea environments, abrupt variations in salinity across adjacent depth levels may amplify gradients during training. To mitigate potential gradient explosion introduced by convolutional operations alone, we further apply a SimpleGate activation module after convolution:
@@
where the weight $\xi$ is a learnable scaling factor used to balance the relative contributions of the original depth features $\mathbf{Z}_{ssp}^{r}$ and the enhanced features $SG\left(\mathbf{\dot{Z}}_{ssp}^{r}\right)$ in the residual link. Through training, the model automatically adjusts the size of $\xi$: when $\xi$ is large, it emphasizes the local nonlinear changes captured by the depthwise separable convolution and SimpleGate; when $\xi$ is small or even close to 0, it focuses more on preserving the overall trend of the SSP.
@@
By synergistically integrating the LFB and FMB, the module aligns multimodal labels with deep spatial context features and enhances them concurrently. This process yields rich, structured, and physically meaningful initial features for MDF-RAGAN, establishing a foundation for robust and precise adversarial learning.
II-B2 Cross-Modal Perturbation Attention Block